OpenStudy (yrelhan4):

In how many different ways three persons A, B, C having 6, 7 and 8 one rupee coins respectively can donate Rs.10 collectively?

OpenStudy (yrelhan4):

@shubhamsrg

terenzreignz (terenzreignz):

Can any one of them choose not to give anything?

OpenStudy (anonymous):

can some give more and some give less??

OpenStudy (yrelhan4):

@sambhav__jain can give max 6, b can give max 7, c can give max 8. i think so @terenzreignz

terenzreignz (terenzreignz):

Then, let's assume A gives 0. The problem becomes how many ways B and C can give 10.

OpenStudy (yrelhan4):

are you sure you can assume a gives nothing?

terenzreignz (terenzreignz):

We can't. But we can consider all the cases where A gives 0, A gives 1, etc.

OpenStudy (yrelhan4):

well that would be quite long. but how would you solve if b and c give10?

terenzreignz (terenzreignz):

Well, there are six ways, depending on how many B gives. B gives 7, C gives 3 B gives 6, C gives 4 B gives 5, C gives 5 B gives 4, C gives 6 B gives 3, C gives 7 B gives 2, C gives 8 B gives 1....C can't give enough... B can't give just 1. Six possibilities if A gives 0 Now check if A gives 1. The problem is how many ways B and C can give 9

OpenStudy (yrelhan4):

hmmm. got it. but this would be too long for an objective question.

OpenStudy (shubhamsrg):

firstly you may wanna read this: http://www.cuemath.com/2012/permutations-and-combinations-counting-the-number-of-solutions-of-an-integral-equation/ let A give x, B gives y and C gives z then x+y+z = 10 where x<=6 y<=7 z<=8 --> 6-x >= 0 7-y >=0 8-z >=0 our expression is x+y+z=10 , we have to find its integer solutions. => (6-x) + (7-y) + (8-z) = (6+7+8)-10 (6-x) + (7-y) + (8-z) = 11 hence our ans will be directly C(11+3-1 ,3-1) = C(13,2) = 78 ways Hope that helped ?

OpenStudy (yrelhan4):

yes i did that too. but the answer is 47. i found this solution... Suppose they are A B C now A+B+C=10 => So the number of integral solution =12C2 Now A cannot give more than 6 so the cases where A gives 7 or more need to be deducted A=A'+7 => So A'+B+C=3 the number of ways = 5C2. => Same way for B it's 4C2 and for C it's 3C2 => Total = 12C2 - (5C2 + 4C2 + 3C2) = 47 why did they do A=A'+7?

OpenStudy (shubhamsrg):

Sorry had got dc. Well yes there is a slight flaw in my method. (6-x) + (7-y) + (8-z) = 11 The formula is incorrectly applied here as it does not acomodate for 6-x>=0, according to it, it treats it as X and this will include X=7, even though that'll make x negative hmm we need to subtract cases from 78 1st case : x>=6 --> 6C2 2nd case: y>=7 --> 5C2 and 3rd case : 4C2 final ans= 78 - 6C2 -5C2 - 4C2 = 47

OpenStudy (shubhamsrg):

am sorry that'll be for x>=7, y>=8 and z>=9

OpenStudy (yrelhan4):

i am sorry. "The formula is incorrectly applied here as it does not acomodate for 6-x>=0, according to it, it treats it as X and this will include X=7, even though that'll make x negative" didnt get that. :O

OpenStudy (yrelhan4):

@shubhamsrg

OpenStudy (yrelhan4):

@Mashy

OpenStudy (anonymous):

combinations ahhh! :P

OpenStudy (yrelhan4):

not your forte? :P

OpenStudy (anonymous):

not really.. but i know the concepts.. but m not getting a lead to start it :P

OpenStudy (yrelhan4):

all that i dont understand here is this... "The formula is incorrectly applied here as it does not acomodate for 6-x>=0, according to it, it treats it as X and this will include X=7, even though that'll make x negative" i dont understand what 'shubhang' meant. :P

OpenStudy (anonymous):

can we do it this way case 1.. B and C give 4 rupee together.. so A has to give 6 work out how many ways this is possible case 2 .. B and C give 5 rupee togther.. so A has to give 5 .. last case.. B and C give 9 rupee togther so A has to give 1 :D

OpenStudy (yrelhan4):

you want me to count 47 ways like that? :P

OpenStudy (anonymous):

no.. only 6 cases like that :-/ see case 1.. you have to find in how many ways can B and C give 4 rupee together :D

OpenStudy (anonymous):

let me try and work out!

OpenStudy (yrelhan4):

hmm. thats pretty much what terenz said. i like shubhams method, but i cant understand it. :P

OpenStudy (anonymous):

so case 1.. B and C can do it as 1,3 2,2 3,1.... hence case 1 is 3 case 2.. B anc C give 5 rupee together.. so 1,4... 2,3...3,2....4,1... hence case 2 is 4 case 3......................6...... togther so 1,5....2,4...3,3...4,2....5,1.. hence case 3 is 5 case 4 ... is similary 6 case 5 is 7 case 6 8 so the answer is 8+7+6+5+4+3 and thats 33.. where did i go wrong? :O

OpenStudy (anonymous):

are we considering zero as well?? not giving as well?

OpenStudy (yrelhan4):

yup.

OpenStudy (anonymous):

ohh.. then Case 1 .. 5 case 2 ..6 case 3.. 7 case 4..8 case 5..9 case 6..10 and one more case (A =0) .. 11 wait now m getting more how? :O .. i just added 2 to each case :-/

OpenStudy (anonymous):

oh damn.. forget it.. i suck at this!

OpenStudy (yrelhan4):

haha. i have to go now. i'll come back and have a look. :P

OpenStudy (anonymous):

me going too bye :P

OpenStudy (anonymous):

@yrelhan4 ,, terenzreignz has a point there

OpenStudy (anonymous):

it leads to the solution

OpenStudy (yrelhan4):

sure he has. yes it leads to the solution. but objective questions generally dont have long solutions.

OpenStudy (shubhamsrg):

Will anyone tell me from where did "shubhang" actually came into existence? -_- "The formula is incorrectly applied here as it does not accommodate for 6-x>=0, according to it, it treats it as X and this will include X=7, even though that'll make x negative" We have to deal with (6-x) + (7-y) + (8-z) = 11 or X + Y + Z =11 where X=6-x and Y=7-y and Z=8-z the formula works for X+Y+Z = 11 , with eqch >=0, i.e. the formulla will also take in cases where X=7 and above but X can't be 7,it can be max 6, if its 7, it'll make x -ve. Sorry my eng is not very good :/

OpenStudy (yrelhan4):

hmm. seems i got it. thank you. and i have no idea where shubhang came from. ask @Mashy ! :P

OpenStudy (shubhamsrg):

hmm, glad to help.